The population of Cartown, USA has decreased by 1/4 over the last year due to an auto plant closing. The population now is 58,215. What was the population last year?
A.14,554
B.46,572
C.77,620

Answer by checkley77(12844)

x-x/4=58,215
(4x-x)/4=58,215
3x/4=58,215
3x=58,215*4
3x=232,860
x=232,860/3
x=77,620 was the population last year.
ANSWER C).